# Recurrent Deep Kernel Learning of Dynamical Systems

### v1

@@ 4 Numerical Results
From high-dimensional and noisy measurements $\mathbf{x}$ of these systems, in our experiments, we aim to (a) denoise of the measurements, (b) learn compact and interpretable representations, (c) predict the dynamics, and (d) quantify uncertainties.
@@
Learning Compact and Interpretable Representations. The second aspect we are interested in is assessing the ability of the framework to learn compact and interpretable latent representations of the system states. To visualize the latent (state) representations, we employ the t-distrubuted stochastic neighbour embedding (t-SNE) method [54] to project the latent states to a 2-dimensional space and inspect their correlation with the true state variables (see Figure 6 and 10).
@@
Quantifying Uncertainties. Ultimately, we would like to quantify uncertainties, deriving from the noisy measurements, properly. However, uncertainties over the latent variables are hard to visualize and, consequently analyze [41]. Therefore, we study uncertainties in the measurement space that be more informative then uncertainties over the latent variables. We generate multiple rollouts of the ROM for a fixed initial condition, we decode the latent-space trajectories back to the measurement space using the decoder, we compute standard deviation of the different trajectories, and we plot heatmaps representing the evolution of the uncertainties over time (see Figure 9 and 13).
@@ 4.1 Double Pendulum
The first example we consider is an actuated double pendulum. The double pendulum exhibits chaotic and nonlinear behaviour, making the prediction of its dynamics from high-dimensional and noisy measurements extremely challenging. Its dynamics can be described as a function of its joint angles $(\theta_{1},\theta_{2})$, velocities $(\dot{\theta}_{1},\dot{\theta}_{2})$, and accelerations $(\ddot{\theta}_{1},\ddot{\theta}_{2})$. We indicate with $\theta_{1},\theta_{2}$ the angle of the first joint and second joint, $l_{1},l_{2}$ the length of the two links, and $m_{1},m_{2}$ the mass of the two links. The equations of motions of the double pendulum can be written as in state-space form as:
@@
To further understand the effect of the recurrent SVDKL architecture on the prediction performance of the ROM, we quantitatively analyse, using the PSNR and L1 metric , the quality of the reconstructions $\hat{\mathbf{x}}_{t}$ and $\hat{\mathbf{x}}_{t+1}$. The results for different values of $H$, where $H=1$ corresponds to the original architecture proposed in [41], and noise variance $\sigma^{2}_{\mathbf{x}}$ are reported in Table 1. Increasing the history length $H$ improves the encoding and latent model performance and consequently denoising abilities of the model. However, due to the sequential nature of the LSTM, longer input sequences requires more computations and may slow down the training. We found that a value of $H\in[10,20]$ provides a good trade-off between results and computational burden.
@@
Figure 7: t-SNE visualisation of the mean of the latent state distribution for different levels of noise $\sigma^{2}_{\mathbf{x}}\in\{0,0.25^{2},0.5^{2}\}$ on the measurements, Figure a), b), and c), respectively. The first angle $\theta_{1}$ is fixed within a range, while the colour bar represents the second angle $\theta_{2}$.
@@
The reconstructions remain close to the true measurements, even with the highest noise level ($\sigma^{2}_{\mathbf{x}}=0.5^{2}$) for the first 20 timesteps. Afterwards, especially in the noisy-measurement cases, we notice a gradual divergence. However, it is worth mentioning that the double pendulum is a chaotic systems and small perturbations of the initial conditions generate drastically different trajectories. Therefore, it is natural to observe this behaviour, especially with additive noise on the measurements acting as a perturbation of the initial conditions.
@@ 4.2 Nonlinear Reaction-Diffusion Problem
where $u=u(x,y,t)$ and $v=v(x,y,t)$ describe the evolution of the spiral wave over time in the spatial domain $(x,y)\in[-10,10]$, and $\beta$ and $d$ are the parameters regulating the reaction and diffusion behaviour of the system, respectively. Our parameter of interest is $\beta$, varying in the range $[0.5,1.5]$. Similarly to [55], we assume periodic boundary conditions and initial condition equal to:
@@
Similarly to the case of the double pendulum, we quantitatively analyse, using the PSNR and L1 metrics, the quality of the reconstructions $\hat{\mathbf{x}}_{t}$ and $\hat{\mathbf{x}}_{t+1}$. The results for different values of $H$, where $H=1$ corresponds to the original architecture proposed in [41], and noise $\sigma^{2}_{\mathbf{x}}$ are reported in Table 2. Even in the reaction-diffusion PDE, increasing the history length $H$ improves the encoding and latent model performance and consequently denoising abilities of the model.
@@
Figure 11: t-SNE visualisation of the mean of the latent state distribution for different levels of noise $\sigma^{2}_{\mathbf{x}}\in\{0,0.25^{2},0.5^{2}\}$ on the measurements, Figure a), b), and c), respectively. The colorbar represents the timestep.
@@
Again, it is possible to notice that the model is sensitive to the noise added to the measurements, as the standard deviation of the predicted trajectory grows with more noise, i.e., the pendulum position is more blurred in the case of $\sigma^{2}_{\mathbf{x}}=0.5^{2}$ than in the case of $\sigma^{2}_{\mathbf{x}}=0.0$.
5 Conclusion

### v3

@@ 4 Numerical Results
We assume to have access only to high-dimensional and noisy measurements $\mathbf{x}$ of these systems. In our experiments, we aim to (a) denoise of the measurements, (b) learn compact and interpretable representations, (c) predict the systems’ dynamics, and (d) quantify uncertainties.
@@
Learning Compact and Interpretable Representations. The second aspect we are interested in is assessing the ability of the framework to learn compact and interpretable latent representations of the system states. To visualize the latent (state) representations, we employ the t-distributed stochastic neighbor embedding (t-SNE) method [60] to project the latent states to a 2-dimensional space and inspect their correlation with the true state variables (see Figure 6 and 10).
@@
Quantifying Uncertainties. Ultimately, we would like to quantify uncertainties, deriving from the noisy measurements, properly. However, uncertainties over the latent variables are hard to visualize and, consequently analyze [41]. Therefore, we study uncertainties in the measurement space that can be more informative then uncertainties over the latent variables. We generate multiple rollouts of the ROM for a fixed initial condition, we decode the latent-space trajectories back to the measurement space using the decoder, we compute standard deviation of the different trajectories, and we plot heatmaps representing the evolution of the uncertainties over time (see Figure 9 and 13).
@@ 4.1 Double Pendulum
The first example we consider is an actuated double pendulum. The double pendulum exhibits chaotic and nonlinear behavior, making the prediction of its dynamics from high-dimensional and noisy measurements extremely challenging. Its dynamics can be described as a function of its joint angles $(\theta_{1},\theta_{2})$, velocities $(\dot{\theta}_{1},\dot{\theta}_{2})$, and accelerations $(\ddot{\theta}_{1},\ddot{\theta}_{2})$. We indicate with $\theta_{1},\theta_{2}$ the angle of the first joint and second joint, $l_{1},l_{2}$ the length of the two links, and $m_{1},m_{2}$ the mass of the two links. The equations of motions of the double pendulum can be written as in state-space form as:
@@
To further understand the effect of the recurrent SVDKL architecture on the prediction performance of the ROM, we quantitatively analyze, using the PSNR and L1 metric , the quality of the reconstructions $\hat{\mathbf{x}}_{t}$ and $\hat{\mathbf{x}}_{t+1}$. The results for different values of $H$, where $H=1$ corresponds to the original architecture proposed in [41], and noise variance $\sigma^{2}_{\mathbf{x}}$ are reported in Table 1. Increasing the history length $H$ improves the encoding and latent model performance and consequently denoising abilities of the model, as shown by a higher PSNR and a lower $L_{1}$ norm. However, due to the sequential nature of the LSTM, longer input sequences requires more computations and may slow down the training. We found that a value of $H\in[10,20]$ provides a good trade-off between results and computational burden.
@@
Figure 7: t-SNE visualization of the mean of the latent state distribution for different levels of noise $\sigma^{2}_{\mathbf{x}}\in\{0,0.25^{2},0.5^{2}\}$ on the measurements, Figure a), b), and c), respectively. The first angle $\theta_{1}$ is fixed within a range, while the color bar represents the second angle $\theta_{2}$.
@@
The reconstructions remain close to the true measurements, even with the highest noise level ($\sigma^{2}_{\mathbf{x}}=0.5^{2}$) for the first 20 timesteps. Afterwards, especially in the noisy-measurement cases, we notice a gradual divergence. However, it is worth mentioning that the double pendulum is a chaotic systems and small perturbations of the initial conditions generate drastically different trajectories. Therefore, it is natural to observe this behavior, especially with additive noise on the measurements acting as a perturbation of the initial conditions.
@@ 4.2 Nonlinear Reaction-Diffusion Problem
where $u=u(x,y,t)$ and $v=v(x,y,t)$ describe the evolution of the spiral wave over time in the spatial domain $(x,y)\in[-10,10]$, and $\beta$ and $d$ are the parameters regulating the reaction and diffusion behavior of the system, respectively. Our parameter of interest is $\beta$, varying in the range $[0.5,1.5]$. Similarly to [61], we assume periodic boundary conditions and initial condition equal to:
@@
Similarly to the case of the double pendulum, we quantitatively analyze, using the PSNR and L1 metrics, the quality of the reconstructions $\hat{\mathbf{x}}_{t}$ and $\hat{\mathbf{x}}_{t+1}$. The results for different values of $H$, where $H=1$ corresponds to the original architecture proposed in [41], and noise $\sigma^{2}_{\mathbf{x}}$ are reported in Table 2. Even in the reaction-diffusion PDE, increasing the history length $H$ improves the encoding and latent model performance and consequently denoising abilities of the model.
@@
Figure 11: t-SNE visualization of the mean of the latent state distribution for different levels of noise $\sigma^{2}_{\mathbf{x}}\in\{0,0.25^{2},0.5^{2}\}$ on the measurements, Figure a), b), and c), respectively. The colorbar represents the timesteps.
@@
Again, it is possible to notice that the model is sensitive to the noise added to the measurements, as the standard deviation of the predicted trajectory grows with more noise, i.e., the spiral is more blurred in the case of $\sigma^{2}_{\mathbf{x}}=0.5^{2}$ than in the case of $\sigma^{2}_{\mathbf{x}}=0.0$.
5 Conclusion and Discussion
@@
In recent years, model and dimensionality reduction [23, 24, 25, 27, 28, 36, 37, 38], uncertainty quantification [5, 6, 7, 26], and measurement denoising [40, 42] have been essential aspects of the research in scientific machine learning. However, novel methods have been very often tailored for only one or two of these challenges at a time. Simultaneously tackling all these challenges is very difficult. To the best of our knowledge, the proposed method in this work is the first one capable of denoising high-dimensional measurements, reducing their dimensionality into interpretable latent spaces, predicting system evolution, and quantifying modeling uncertainties simultaneously, as shown in Section 4.